OpenStudy (anonymous):

How do I convert 25.875 into feet.???

OpenStudy (juicstice):

what is the unit used for 25.875?

OpenStudy (triciaal):

you are missing the units in the original question @Juicstice is waiting for a response

OpenStudy (anonymous):

@rtory78 there must suppose to have two units.example How do i convert 25.875 inch into feet?

OpenStudy (freckles):

pretend we have 25.875 W where W is a unit I made up. and assume t is some number. And pretend t W=a feet so we can convert 25.875 W to feet by doing? \[25.875 W=25.875 W \cdot \frac{a ft}{t W} \\ \text{ we see the W units cancel out and we have: } \\ 25.875 =25.875 \cdot \frac{a ft}{t} =\frac{25.875 a}{t} ft\]

Directrix (directrix):

Whatever the conversion, if the units are of the same type such as length, use the online conversion tool. If your question were the following: How do I convert 25.875 METERS into feet? Go to Google and enter meters to feet. A conversion tool pops up and you may enter the type conversion you want as well as the associated numbers. @rtory78